The information below relates to a closed economy without a government. use the information to answer the questions that follows. Autonomous spending R285 million, Autonomous investment R300 million, marginal propensity to consume 0.4.
1. Calculate the level of autonomous spending.
2. Calculate the size of the multiplier.
3. Calculate the equilibrium level of income.
4, Suppose the equilibrium level of income is 1000 and it is below the full employment level of income of 1200. Calculate the change in investment spending required to reach full employment.
1
Expert's answer
2020-09-30T08:59:54-0400
  1. Autonomous Spending "=\bar C + \bar I = 285+300=585"
  2. Multiplier "= \frac{1}{1-MPC} = \frac{1}{1-0.4} = \frac{1}{0.6} = 1.66666667"
  3. Equilibrium Level of Income "Y=C+(MPC)Y+I =285+0.4Y+300
\to 0.6 Y = 585 \to Y=\frac{585}{0.6} = 975"
  4. Investment spending required to reach full employment equilibrium

"1200= C+ (MPC)1000 + I"

"1200=285+400+I"

"1200=685+I"

"I=515"